Three times the difference if 5 minus a number is 27. What's the number
Answer by addingup(3677)

3(x-5)= 27
3x-15= 27
3x= 42
x= 14
Proof, put the 14 in the original problem and see what happens:
3(14-5)= 27
42-15= 27
27= 27 We have an equality, we have the right answer.